OpenStudy (anonymous):

can sme on plzz help mee Solve using substitution: 9y − x − 8 = 0 -3x + 7y = -16 Answer (-8, 0) (1, 1) (0, 8/9) (10, 2)

OpenStudy (anonymous):

cn yu hlp me plzz alexeis_nicole

OpenStudy (anonymous):

so is it asking you to chose one set of coordinate points that apply to both of the equations?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay so lets work with the first one.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

(-8,0) = (x,y) so if we substitute the x value with -8 and the y with 0 in the first equation .... it'll look like this 9(0)-(-8)-8 = 0 0+8-8=0 0=0 So for this equation, the coordinate (-8,0) works do you wanna try this coordinate for the next equation? :)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay do you knw how to do it?

OpenStudy (anonymous):

I think I do

OpenStudy (anonymous):

okay show me :)

OpenStudy (texaschic101):

you could just solve the system of equations with substitution by getting x by itself, then you would not have to go through every answer choice

OpenStudy (anonymous):

-3x + 7y = -16 -3(0)+7(-8)-16 0-56-16 that's all I kn ii think its wrong

OpenStudy (anonymous):

wait. you dont have to bring -16 to the other side. -3(0) -56 = -16 0-56=-16 -56=-16 so although (-8,0) satisfies the first eqn it does not satisfy the second

OpenStudy (texaschic101):

-x = 8 - 9y (multiply by -1) x = 9y - 8 now just sub 9y - 8 in for x in the other equation and solve for y. Once y is found, sub that back into the first equation and solve for x. You do not have to go through every answer choice.

OpenStudy (anonymous):

okay can we do sme mre well if yu wnt to

OpenStudy (anonymous):

or do what texaschic101 said

OpenStudy (anonymous):

uhmm okay

OpenStudy (anonymous):

well it's up to you whatever works, i dont mind going through every coordinate with you

OpenStudy (anonymous):

Identify the system of inequalities that would generate the following graph: Answer y < 7x − 3; y > 3x − 2 y < 7x − 3; y ≥ 3x − 2 y > 3x − 7; y < 2x − 3 y < 3x − 7; y ≤ 2x − 3 want to kn hw to do this

OpenStudy (anonymous):

where's the graph

OpenStudy (anonymous):

?

OpenStudy (anonymous):

I mean this Select an ordered pair from the choices below that is a solution to the following system of inequalities: 3x < -3+ 2y y − 2x > 4 Answer (0, 4) (-6, 0) (0, 6) (6, 3

OpenStudy (anonymous):

?

OpenStudy (anonymous):

okay... so substitute the y-values and x-values

OpenStudy (anonymous):

do (0,6)

OpenStudy (anonymous):

okay the same way we been doing it